by weight one fertilizer is 20% potassium, 30% nitrogen and 50% phosphorus. a second fertilizer has percents of 10%, 20% and 70% respectively and a third fertilizer has percentage of 0%, 30% and 70% respectively. how much of each should be mixed to get 200 kg of fertilizers with percents of 12%, 25% and 63% respectively?
Answer by Theo(13342)

x = amount of fertilizer 1
y = amount of fertilizer 2
z = amount of fertilizer 3

it helps to make a table for this.

your table would look like this:

amount of fertilizer in kilograms
fertilizer     total     potassium     nitrogen     phosphorus

1                x          .2x          .3x           .5x
2                y          .1y          .2y           .7y
3                z          ---          .3z           .7z 
4              200           24           50           126

the amounts of potassium, nitrogen, and phosphorous for fertilizer 4 are given to you by weight because you are told the proportions and you know what the total weight is.

potassium = .12 * 200 = 24
nitrogen = .25 * 200 = 50
phosphorus = .63 * 200 = 126

you wind up with 3 equations that need to be solved simultaneously.

note that these equations are from the vertical columns for potassium, nitrogen, and phosphorus in your original table above.

those equations are:
           .2x + .1y + .0z = 24  (equation 1)
           .3x + .2y + .3z = 50  (equation 2)
           .5x + .7y + .7z = 126 (equation 3)

once you solve these simultaneously, you will have the amounts of x,y,z that you need which are the amounts of fertilizer 1, fertilizer 2, fertilizer 3 that you need.

we will eliminate the decimals first so they won't bother us.

multiply all equations by 10 to get:

           2x + 1y + 0z = 240  (equation 1)
           3x + 2y + 3z = 500  (equation 2)
           5x + 7y + 7z = 1260 (equation 3)

first we will take equation 1 and 2 and eliminate the y variable.

we will do that by multiplying equation 1 by 2 and then subtract equation 2 from equation 1 to get equation 4.

           4x + 2y + 0z =  480  (equation 1 multiplied by 2)
 minus:    3x + 2y + 3z =  500  (equation 2)
 equals:   1x      - 3z =  -20  (equation 4)

next we will take equation 2 and 3 and eliminate the y variable.

we will do that by multiplying equation 2 by 7 and multiplying equation 3 by 2 and then subtracting equation 3 from equation 2.

           21x + 14y + 21z = 3500  (equation 2 multiplied by 7)
 minus:    10x + 14y + 14z = 2520  (equation 3 multiplied by 2)
 equals:   11x       +  7z = 980   (equation 5)

next we will take equation 4 and equation 5 and eliminate the x variable.

we will do that by multiplying equation 4 by 11 and then subtracting equation 5 from equation 4

                  11x - 33z =  -220   (equation 4 multiplied by 11)
      minus:      11x +  7z =   980   (equation 5)
      equals:         - 40z = -1200


we solve for z to get:

z = 30

we use this value of z and substitute in equation 5 to get:
11x + 7z = 980 becomes:
11x + 7*(30) = 980 which becomes:
11x + 210 = 980
subtract 210 from both sides to get:
11x = 980-210 = 770
divide both sides by 11 to get:
x = 70

next we use the values of z = 30 and x = 70 and go to equation 3 to get:
5x + 7y + 7z = 1260 becomes:
5*(70) + 7y + 7*30) = 1260 which becomes:
350 + 7y + 210 = 1260 which becomes
7y + 560 = 1260
subtract 560 from both sides to get:
7y = 1260 - 560 = 700
divide both sides by 7 to get:
y = 100

we have values for x,y,z.

they are:

x = 70
y = 100
z = 30

we need to test these values to see if they are good.

our original equations are:

           .2x + .1y + .0z = 24  (equation 1)
           .3x + .2y + .3z = 50  (equation 2)
           .5x + .7y + .7z = 126 (equation 3)

first we check to see if x + y + z = 200.

it does.

that's good because the total of fertilizer 1, 2, 3 should equal 200 and it does.

next we substitute in all 3 equations to get:

           .2x + .1y + .0z = 24  (equation 1)
           .3x + .2y + .3z = 50  (equation 2)
           .5x + .7y + .7z = 126 (equation 3)

becomes:
           .2*70 + .1*100 + .0*30 = 24  (equation 1)
           .3*70 + .2*100 + .3*30 = 50  (equation 2)
           .5*70 + .7*100 + .7*30 = 126 (equation 3)

we simplify to get:
           14 +  10 +  0 =  24  (equation 1)
           21 +  20 +  9 =  50  (equation 2)
           35 +  70 + 21 = 126  (equation 3)
           -------------------
  total:   70 + 100 + 30 = 200  (sum of equations 1,2,3)

Looks like we have the answer.

he needs:
70 pounds of fertilizer 1
100 pounds of fertilizer 2
30 pounds of fertilizer 3

he will get:

200 pounds of fertilizer 4 that contains:

24 pounds of potassium
50 pounds of nitrogen
126 pounds of phosphorus